# Real-time Rendering-based Surgical Instrument Tracking via Evolutionary Optimization

### 12 Mar 2026 v1

@@ V Discussions and Conclusions
A remaining limitation of rendering-based methods lies in their dependence on accurate segmentation. Errors caused by occlusion or noisy segmentation propagate directly to pose and kinematic estimation. Furthermore, the SurgPose dataset does not contain sequences with overlapping instruments, preventing evaluation of our bi-manual tracking method under frequent instrument–instrument occlusions encountered in real surgical procedures. In future work, we will incorporate more robust instrument-level segmentation strategies to better handle occlusion scenarios. We also plan to leverage high-throughput batch rendering platforms (e.g., [10.1145/3680528.3687629]) to fully exploit the blackbox optimization paradigm of evolutionary strategies and further improve computational scalability.

### 13 Mar 2026 v2

@@ V Discussions and Conclusions
A remaining limitation of rendering-based methods lies in their dependence on accurate segmentation. Errors caused by occlusion or noisy segmentation propagate directly to pose and kinematic estimation. Furthermore, the SurgPose dataset does not contain sequences with overlapping instruments, preventing evaluation of our bi-manual tracking method under frequent instrument–instrument occlusions encountered in real surgical procedures. In future work, we will incorporate more robust instrument-level segmentation strategies to better handle occlusion scenarios. We also plan to leverage high-throughput batch rendering platforms (e.g., [37]) to fully exploit the blackbox optimization paradigm of evolutionary strategies and further improve computational scalability.